# Kinematic Tokenization: Optimization-Based Continuous-Time Tokens for Learnable Decision Policies in Noisy Time Series

### v1

@@ 6.1 Signal Deficits and Liquidation Equilibrium
Model Action Distribution (NVDA, $N=760$) Performance
@@
PatchTST 0 760 0 100.0% 0.0% N/A
SplineGPT 303 229 228 70.0% +581.6% 2.33
Table 7: The Liquidation Equilibrium of Discrete Tokens. We diagnosed the specific action distribution of the baseline (PatchTST) versus SplineGPT at the standard volatility threshold ($\tau=1.0\%$). While SplineGPT maintains a balanced trading policy, the discrete baseline collapses to a monotonic “Always Sell” policy. This indicates that the discrete representation fails to generate sufficient signal confidence to overcome the asymmetric risk penalty ($w_{sell}=10.0$), forcing the model into the safest possible state (100% Cash).
@@ 6.1.1 Sensitivity Sweep Analysis
To verify that this failure was not merely a result of an overly conservative threshold, we performed a sensitivity sweep, lowering the volatility threshold to $\tau=0.5\%$ to encourage model activity. As shown in Table 8, the discrete baseline failed to recover even in this aggressive regime, remaining trapped in the Liquidation Equilibrium. Conversely, SplineGPT adapted successfully, increasing its action rate to $\approx 82\%$ while maintaining a robust Sharpe ratio ($1.42$). This provides evidence that the gap is driven primarily by representation (and not only by threshold conservatism), since PatchTST remains in the Liquidation Equilibrium even when $\tau$ is reduced.
@@
Conservative 100.0% (Sell) N/A 35.4% 1.82 Selective Profitable
($\tau=1.0\%$)
Aggressive 100.0% (Sell) N/A 82.1% 1.42 Active Profitable
($\tau=0.5\%$)
Table 8: Behavioral Divergence Under Sensitivity Sweep. We lowered the volatility threshold ($\tau$) from 1.0% to 0.5% to force baselines to trade. While SplineGPT maintained robust risk-adjusted returns (Sharpe >1.0>1.0), PatchTST remained stuck in the Liquidation Equilibrium (100% Sell / 0.0% Return), failing to identify actionable entry points even when the barrier to entry was halved.
6.2 Explicit vs. Implicit Smoothing (PatchTST)
@@
6.3 Mechanisms of Alpha Generation
Given the observed separation between SplineGPT and the abstaining baselines, we highlight two recurring regimes in which the learned policy appears to benefit from smooth coupled derivatives:
  1. 1.
    Momentum Convexity (TSLA, INTC): In high-beta assets, price action often exhibits significant convexity (acceleration) during breakouts. SplineGPT successfully captured the positive ”Jerk” ($c_{3}$) associated with the onset of these rallies. By acting on the smooth derivative rather than the noisy price difference, the model entered positions earlier and held through minor volatility that shook out the discrete baselines.
  2. 2.
    Crash Avoidance (PFE): The most critical validation of the ”Active Risk Management” objective is observed in the Pfizer (PFE) case. While the asset shed over 40% of its value, the model preserved capital (MaxDD -18.8%). As visualized in Figure 7, the kinematic signatures during this crash are complex—characterized by volatility clustering in the higher-order derivatives. The model’s ability to navigate this crash demonstrates its capacity to decouple meaningful structural breaks from transient noise within the 9-dimensional manifold.
@@ 6.4 Objective-Induced Conservatism
In parabolic regimes (e.g., NVDA, JPM), SplineGPT yields positive returns but trails the Buy & Hold benchmark. This divergence is not a failure of prediction, but a structural consequence of the asymmetric risk objective (Eq. 10). Standard trend-following optimizes for total return ($R$), implicitly accepting infinite drawdown risk. In contrast, our objective functions as a constrained optimization problem: maximizing $R$ subject to strict downside penalties. Consequently, the model exhibits convexity shedding—exiting positions during intermediate mean-reversion events to reset risk exposure. This behavior characterizes the model as a “regime-switching” agent rather than a pure momentum factor, trading absolute ceiling for a higher risk-adjusted floor (Sharpe/Sortino).
@@ 6.5 Sensitivity to Execution Costs & Turnover
We observe high annualized turnover (36x–86x), indicating that the model has learned a short-horizon active trading policy rather than a passive holding strategy. Despite this high-turnover behavior, the strategy demonstrates significant resilience to friction:
@@
INTC 0.52 0.44 0.36 0.19 75.2x
NVDA 1.17 1.08 0.99 0.81 74.5x
JPM 0.94 0.86 0.77 0.59 36.8x
XOM 0.23 0.11 -0.03 -0.31 47.7x
PFE -0.62 -0.74 -0.87 -1.13 45.9x
TSLA 0.84 0.75 0.67 0.49 86.5x
Table 9: Execution Sensitivity. We report Sharpe Ratios under varying transaction cost assumptions. The strategy maintains positive expectancy for momentum-driven assets up to 20bps, though performance degrades for mean-reverting or low-volatility names (e.g., XOM, PFE) as costs increase. The high turnover values confirm the model operates as a high-frequency daily tactical agent.

### v2

@@ 6.1 Signal Deficits and Liquidation Equilibrium
Model Action Distribution Performance
@@
PatchTST 0 762 0 100.0% 0.0% N/A
SplineGPT 294 228 240 68.5% +579.2% 2.31
Table 5: The Liquidation Equilibrium of Discrete Tokens. We diagnosed the specific action distribution of the baseline (PatchTST) versus SplineGPT at the standard volatility threshold ($\tau=1.0\%$) for NVDA. While SplineGPT maintains a balanced trading policy, the discrete baseline collapses to a monotonic “Always Sell” policy. This indicates that the discrete representation fails to generate sufficient signal confidence to overcome the asymmetric risk penalty ($w_{sell}=10.0$), forcing the model into the safest possible state (100% Cash).
@@ 6.1.1 Sensitivity Sweep Analysis
To verify that the baseline failure was not merely a result of an overly conservative threshold, we performed a sensitivity sweep, lowering the volatility threshold to $\tau=0.5\%$ to encourage model activity. As shown in Table 6, the discrete baseline failed to recover even in this aggressive regime, remaining trapped in the Liquidation Equilibrium. Conversely, SplineGPT adapted successfully, increasing its action rate to $90.8\%$ while improving its Sharpe ratio to $2.48$ and better capturing the extreme bullish nature of the underlying NVDA equity by being actively invested more often. This demonstrates that the performance gap is representational where discrete measurement tokens (versus discrete spline tokens) lack the signal-to-noise ratio required to support non-trivial decision boundaries under this standard PatchTST architecture, regardless of the threshold.
@@
Very Aggressive 100.0% (Sell) N/A 100.0% 2.73 Active High-Yield
($\tau=0.25\%$)
Aggressive 90.8% 2.48 Active High-Yield
($\tau=0.50\%$)
Conservative 67.1% 2.20 Selective Robust
($\tau=1.00\%$)
Very Conservative 38.3% 1.74 Precision Defensive
($\tau=2.00\%$)
Table 6: Behavioral Divergence Under Sensitivity Sweep (NVDA). We varied the volatility threshold ($\tau$) from 0.25% to 2.0%. While SplineGPT scaled its activity rationally—from 100% activity in aggressive regimes to 38% precision in conservative ones—the discrete baseline (PatchTST) remained trapped in the Liquidation Equilibrium (100% Sell) across all thresholds. In the most aggressive regime SplineGPT also exhibits a 100% action rate, but this includes a mix of both buy and sell decisions. This confirms that the baseline’s failure is a fundamental signal deficit, not a tuning artifact.
6.2 Explicit vs. Implicit Smoothing
@@ 6.3 Objective-Induced Conservatism
In parabolic regimes (e.g., NVDA, JPM), SplineGPT yields positive returns but trails the Buy & Hold benchmark. This divergence is not a failure of prediction, but a structural consequence of the asymmetric risk objective (10). Standard trend-following optimizes for total return ($R$), implicitly accepting infinite drawdown risk. In contrast, our objective functions as a constrained optimization problem: maximizing $R$ subject to strict downside penalties. Consequently, the model exhibits convexity shedding—exiting positions during intermediate mean-reversion events to reset risk exposure. This behavior characterizes the model as a “regime-switching” agent rather than a pure momentum factor, trading absolute ceiling for a higher risk-adjusted floor (Sharpe/Sortino).
@@ 6.4 Sensitivity to Execution Costs & Turnover
We observe high annualized turnover (36x–89x), indicating that the model has learned a short-horizon active trading policy rather than a passive holding strategy. Despite this high-turnover behavior, the strategy demonstrates significant resilience to friction:
@@
INTC 1.31 1.19 1.06 0.83 77.7x
NVDA 2.39 2.26 2.12 1.87 70.3x
JPM 1.15 1.07 0.99 0.82 36.0x
XOM 0.73 0.62 0.50 0.27 47.6x
PFE -0.31 -0.43 -0.56 -0.83 44.6x
TSLA 2.05 1.90 1.76 1.48 89.7x
Table 7: Sensitivity to Transaction Costs. We report Sharpe Ratios under varying execution cost assumptions (0–20 basis points per trade). Despite high turnover (>35x>35x), momentum assets (NVDA, TSLA, INTC) retain strong risk-adjusted returns even at 20bps, confirming that the kinematic signal captures substantial price moves that exceed typical friction costs.
6.5 Mechanisms of Alpha Generation
Given the observed separation between SplineGPT and the baselines exhibiting Liquidation Equilibrium, we highlight two recurring regimes in which the learned policy appears to benefit from smooth coupled derivatives:
  1. 1.
    Momentum Convexity (TSLA, INTC): In high-beta assets, price action often exhibits significant convexity (acceleration) during breakouts. SplineGPT successfully captured the positive ”Jerk” ($c_{3}$) associated with the onset of these rallies. By acting on the smooth derivative rather than the noisy price difference, the model entered positions earlier and held through minor volatility that may have induced selling in less sophisticated strategies.
  2. 2.
    Crash Avoidance (PFE): The most critical validation of the ”Active Risk Management” objective is observed in the Pfizer (PFE) case. While the asset shed over 40% of its value, the model preserved capital (MaxDD -18.8%). As visualized in Figure 7, the kinematic signatures during a crash are complex, characterized by volatility clustering in the higher-order derivatives. The model’s ability to navigate this crash demonstrates its capacity to decouple meaningful structural breaks from transient noise within the 9-dimensional manifold.